OpenStudy (anonymous):

g(x) =√x − 3 find (g)g(x)) or (g o g) & the domain

OpenStudy (anonymous):

im having trouble doing the square root of a square root.. and i cant find the domain

OpenStudy (anonymous):

and to clear things up x-3 is all under the radical.. not just x

OpenStudy (anonymous):

\[g(g(x))=\sqrt{\sqrt{x-3}-3}\]

OpenStudy (anonymous):

the root of the root is the fourth root think of exponents: 1/2 of 1/2 is 1/4

OpenStudy (anonymous):

domain. well, we need a number greater than or equal to zero to take a square root. \[\sqrt{x-3}\ge 3\]and,\[x \ge 12\]

OpenStudy (anonymous):

thank you so much esiedl

OpenStudy (anonymous):

np :)